OpenStudy (anonymous):

Prove that ab = 1/2[ (a+b)^2 - (a^2+b^2) ]

OpenStudy (kirbykirby):

Expand the right side of the equation

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

Exactly how? I try but I get stuck after I expand (a+b)^2

OpenStudy (kirbykirby):

\((a+b)^2=a^2+2ab+b^2\) The rest should be straightforward o.O... just remember that if you have a negative infront of parentheses like this: \(-(x+y)\) , the negative get's distributed: \(-(x+y)=-x-y\)